Session: Nov 2001 (session 3)
Wavelength: 6.0 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (C. Bortolotti at Thu Nov 15 23:22:34 MET 2001) Success
No knwn problems.

Effelsberg (A. Kraus at Fri Nov 16 11:05:44 MET 2001) Minor failures
A short break was necessary between UT1217-1223 for a check of two azimuth gears. At UT1515 our telescope control computer failed, and we had to stop. We continued at UT1720, when the backup system was ready.

Torun (K.Borkowski at Fri Nov 16 17:24:10 MET 2001) Failure
Did not participate because of continued terminal problems.

Jodrell_Bank (Alastair Gunn at Mon Nov 19 13:28:21 MET 2001) Success
Tsys measurements wrong for this measurement. Requires measurement of Tcal at lower end of band. Will supply this later.

Cambridge (Alastair Gunn at Mon Nov 19 13:28:45 MET 2001) Success
No known problems.

Urumqi (Liu Xiang at Tue Nov 20 03:43:25 MET 2001) Minor failures
the date in antenna computer might be wrong, it was found after this experiment. I am sorrow, please try to find.

Onsala (John Conway at Wed Nov 21 13:17:29 MET 2001) Minor failures
The first 2/3 of the expt was missed due to high winds. Unfortunately when observations started the phase-cal was left on.

Westerbork (Tony Foley at Thu Nov 22 11:45:49 MET 2001) Success
No known problems